# Dose-aware Diffusion Model for 3D Low-dose PET: Multi-institutional Validation with Reader Study and Real Low-dose Data

### v1

@@ I Introduction
Positron Emission Tomography (PET) is a functional imaging modality widely used in oncology, cardiology, and neurology studies [1, 2, 3]. Given the growing concern of radiation exposure and potentially increased cancer risks accompanied with PET scans, reducing the PET injection dose is desirable [4]. However, PET image quality is negatively affected by the reduced injection dose and may affect diagnostic performance such as the identification of low-contrast lesions [5]. Therefore, reconstructing high-quality images from noisy input is an important topic. In fact, making the radiation dose as low as reasonably achievable (the ALARA principle) is a commonly accepted practice in clinical settings.
Iterative methods such as Maximum Likelihood Expectation Maximization (MLEM) [6] and Ordered Subset Expectation Maximization (OSEM) [7] are commonly used for PET reconstructions. However, they are vulnerable to noise in low-dose PET data.
Deep learning (DL) has emerged as a new reconstruction and post-reconstruction processing approach for medical imaging tasks [8], and many different DL methods were proposed for low-dose PET image reconstructions/restorations [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. However, existing models often tailored to specific hospital and scanner, patient population, acquisition protocol, and image noise level. They typically have limited generalizability across the diverse PET image acquisition settings in reality. Here, we introduced a diffusion-based model [21, 22] to achieve dose-aware 3D PET image denoising.
@@
Recently, diffusion models have become the new state-of-the-art generative models [23, 24]. They are capable of generating high-quality samples from Gaussian noise input, and have demonstrated strong potential for low-dose PET imaging. For example, using the DDPM framework [25], Gong et al. proposed to perform PET image denoising with MRI as prior information for improved image quality [26]. However, PET-MR systems are not widely available and it would be time-consuming to obtain sequential and paired PET-MR images in reality for network training and image reconstructions. Jiang et al.[27] adopted the latent diffusion model [28] for unsupervised PET denoising. Moreover, diffusion models have also been proposed for other imaging modalities, such as CT [29] and MRI [30, 31].
@@
However, these previous diffusion works focus on 2D and do not address the 3D imaging problem. This is particularly important for PET imaging as PET is intrinsically a 3D imaging modality. A 3D diffusion model is desirable, however, due to the hardware memory limit, directly extending the diffusion model to 3D would be difficult. There are a few previous works that aim to address the 3D imaging problems of diffusion models. For example, Chung et al.[32] proposed to apply a TV penalty term along the z-axis to remove inconsistencies within each reverse sampling step in the diffusion model. On the other hand, Lee et al.[33] utilizes 2 pre-trained perpendicular 2D diffusion models to remove inconsistencies between slices. Nonetheless, presented later in this paper, these previous methods failed to produce satisfactory results for 3D low-count PET(e.g, inaccurate organ boundaries in extreme low-dose settings, inconsistent 3D reconstructions, distorted image features in some cases).
@@
Another challenge with PET image denoising is the high variation of image noise levels among patients. The noise level in PET images can be affected by various factors, including: (1) Variations of post tracer injection acquisition start time. (2) Occasional tracer injection infiltration in some patients, causing the tracer stuck in the arm, resulting in high image noise in the body. (3) The variation of patient’s weight and associated weight-based tracer injection protocols. (4) Different hospitals using scanners with varying sensitivities have different total scan times for patients. All these factors result in variability of noise levels in the reconstructed images. Because of this, a method to adaptively denoise images with varying noise levels is desirable. However, previously proposed methods mentioned above have limited generalizability to different noise levels [19]. A network trained on one noise level often fails to produce high-quality reconstructions on other noise levels [19]. To address this problem, we previously proposed to combine multiple U-net-based [34] sub-networks with varying denoising power to generate optimal results for any input noise levels [19]. However, training multiple sub-networks requires tedious data pre-processing and long training time. The testing time also linearly increases with the number of sub-networks.
@@
Moreover, different from other imaging modalities, PET was developed as a quantitative tool, being recognized as providing an objective, and more accurate measure for prognosis and response monitoring purposes than visual inspection alone [35]. However, our experimental results showed that, although standard diffusion models (DDPM [25], DDIM[36]) produce visually appealing images, they typically fail to maintain accurate quantification and produced images with biased tracer uptake. Specifically, the tracer distribution in the entire image change after diffusion models. This issue is more severe in whole-body PET scans since the tracer uptake could vary significantly among organs.
@@
In this work, we developed a dose-aware diffusion model for 3D PET Imaging (DDPET-3D) to address these limitations. The main methodological contributions of the proposed DDPET-3D framework are: (1) We proposed a dose-embedding strategy that allows noise-aware denoising. DDPET-3D can simultaneously denoise 3D PET images with varying low-dose/count levels. (2) We proposed a 2.5D diffusion strategy with multiple fixed noise variables to address the 3D inconsistency issue between slices. DDPET-3D maintains a similar memory burden to 2D diffusion models while achieving high-quality reconstructions. (3) To achieve a quantitative-accurate reconstruction, we proposed to use a denoised prior in DDPET-3D. The proposed denoised prior also allows the diffusion model to converge within 25 sampling steps. DDPET-3D can denoise 3D PET images within a reasonable time constraint ($\sim$15 mins on a single GPU). Previous diffusion methods using DDPM sampling would require approximately $6$ hrs. Further details on network architecture will be provided later in this paper (see Methods).
@@
The ultimate goal of DL methods for PET image denoising is to achieve low-dose PET without compromising image quality and clinical performance. As outlined in Fig. 1(c), evaluated on real-world, large-scale, cross-center, ultra-low-dose 3D PET data, with confirmation from reader studies with nuclear medicine physicians, DDPET-3D demonstrates superior quantitative and qualitative results that are comparable or even better than the 100% full-count data as well as previous baseline methods across varying low-count-levels. DDPET-3D also showed superior generalizability when directly applied to data acquired at a different hospital without further network fine-tunning. We also evaluated the proposed method on a cohort of human studies with real low-dose injection. Presented results show the potential of DDPET-3D to achieve low-dose PET imaging while maintaining image quality.
II Results
@@
[…]
II-A Reader study on three commercial scanners
Figure 3: Sample low-count denoised images reconstructed using the proposed DDPET-3D with different low-count levels. 6 different patients at different low-count levels were randomly chosen from the testing dataset. DDPET-3D can produce consistent denoised results for a wide range of different noise-levels. Ground-truth images were reconstructed using 100% full-count data. Blue arrows point to true-positive lesions or hot spots in different patients. Green arrows point to false-positive hot spots in the images. Blue dashed boxes indicate zoom-in ROI.
We performed a reader study to evaluate the network performance. 45 patient studies acquired using three different commercial scanners from three hospitals (15 from each hospital) were randomly selected from the testing dataset and included for the reader study. Three types of commercial scanners are the Siemens mCT scanner, the Siemens Vision Quadra scanner, and the United Imaging uExplorer scanner. Three hospitals are the Yale New Haven Hospital in the US, Shanghai Ruijin Hospital in China, and the University Hospital of Bern in Switzerland.
Low-count images were reconstructed by down-sampling the PET listmode events. For the images acquired on the Siemens Vision Quadra and the United Imaging uExplorer scanners, six low-count levels were available, including 1%, 2%, 5%, 10%, 25%, and 50%. For the images acquired using the Siemens mCT scanner, three low-count levels were available, including 5%, 10%, and 20%. Images were reconstructed using vendors’ software from United Imaging Healthcare and Siemens Healthineers. Details on image reconstruction is provided later in this paper (see Methods).
All three readers agreed that the proposed DDPET-3D can generate denoised results that are better or at least comparable to the 100% full-count images in terms of overall image quality across three types of commercial scanners at different hospitals. Sample denoised images acquired at the Shanghai Ruijin Hospital (United Imaging uExplorer scanner) are presented in Fig. 3. DDPET-3D can generate consistent denoised results across a wide range of low-count levels. DDPET-3D was able to recover fine details in the images (e.g., rib bone in 1% results; myocardium, spine, and femur in 2% results). DDPET-3D also effectively suppressed background noise (e.g., liver in patients shown in Fig. 3), and lesions became clearer in the denoised images (blue arrows in Fig. 3). DDPET-3D effectively removed false-positive hot spots (green arrows in the 25% results) in the images and the true-positive lesions are easier to identified.
II-B Reader study on real low-dose scans
Figure 5: Sample real low-dose denoised images generated using the proposed DDPET-3D and the UNN[19] methods. Blue arrow points to the aorta wall, which was better reconstructed in DDPET-3D.
II-C Generalizability test in a different hospital
Figure 6: Low-count denoised images reconstructed using the proposed DDPET-3D with different low-count levels for data acquired at the UC Davis Medical Center. The model trained with Shanghai Ruijin data was directly applied for image reconstructions. DDPET-3D can generalize to different medical centers without further fine-tuning.
10 patient studies from the UC Davis Medical Center were included in this experiment. Five low-count levels were generated through listmodel rebinning, including 2.5%, 6.25%, 12.5%, 25%, and 50%. Sample denoised results acquired at the UC Davis Medical Center are presented in Fig. 6. The proposed DDPET-3D showed superior generalizability when directly applied to data acquired at a different hospital. Corresponding quantitative measurements are included in Table I.
II-D Comparison with other baseline denoising methods
Figure 8: Comparison between DDPET-3D and other baseline denoising methods. Blue arrows point to a hot spot that was better recovered by the proposed DDPET-3D. Dahsed blue line indicate the location of the transverse slice. All the images were displayed on the same scale and normalized by the total injected activities.
TABLE I: Quantitative assessment for different methods on all the testing patients from United Imaging and Siemens scanners. The measurements were obtained by averaging the values on the testing human studies. The proposed method consistently produced promising denoised results regardless of input count levels. The best results among different low-count levels are marked in red. Note that 674 out of 774 testing patients from the Shanghai Ruijin Hospital do not have 50% count reconstructions.
[…]
Input 44.929 / 0.474 / 0.831 48.040 / 0.331 / 0.900 51.272 / 0.227 / 0.947
UNN 48.975 / 0.291 / 0.933 50.514 / 0.243 / 0.954 51.970 / 0.205 / 0.969
[…]
TABLE II: Quantitative assessment for different methods. The measurements were obtained by averaging the values on the testing human studies. The proposed method consistently produced promising denoised results regardless of input count levels. $p<0.01$ was observed in all groups when comparing the proposed method with all other methods (except for 50% SSIM values). The best results among different low-count levels are marked in red. Black voxels were removed for calculations.
[…]
The proposed DDPET-3D was compared with UNN [19], DiffusionMBIR [32], and TPDM [33]. As described previously, UNN was chosen because it was also proposed to achieve noise-aware denoising and was one of the winning method in 2022 Ultra Low-dose PET Imaging Challenge. DiffusionMBIR and TPDM were selected because they were recent works proposed to address the diffusion 3D inconsistencies problem in medical imaging. We also compared with the standard DDIM (Denoising Diffusion Implicit Models) sampling [36]. Detailed implementations of different methods will be provided later in this paper (see Methods).
As presented in Fig. 8, standard DDIM sampling produced images with severe inconsistencies between slices in the sagittal view. Both DiffusionMBIR and TPMD improved the results, but the problem still exist. Proposed DDPET-3D method produced noticeably more consistent reconstructions. Moreover, even though other diffusion models produced visually realistic PET images, they failed to accurately reconstruct the structure of different organs. For example, all other diffusion models produced images with distorted myocardium as shown in the transverse slice in Fig. 8. The proposed DDPET-3D method also recovered some fine details in the images, such as the hot spot pointed by the blue arrows in Fig. 8, which is either over-smoothed (in UNN) or barely visible (in other diffusion methods).
Images were quantitatively evaluated using SSIM (Structural Similarity Index), PSNR (Peak signal-to-noise ratio), and RMSE (Root-mean-square error). To facilitate the testing process, we tested all the comparison methods using 20 patients from the entire testing dataset. There were 6 different low-count levels of all the patients, resulting in a total of 120 testing studies. Quantitative results are presented in Table II.
Several ablated experiments were conducted to demonstrate the effectiveness of different proposed components in DDPET-3D. Results from ablation experiments are provided later in this paper (see Ablation studies).
III Discussion and Conclusion
We introduced DDPET-3D, a dose-aware diffusion model for 3D low-dose PET imaging. Evaluated on a large-scale, multi-center, and cross-scanner dataset, low-dose/low-count denoised images generated by DDPET-3D performed favourably or comparably relative to the images reconstructed with 100% full-count. Tested on images acquired using three state-of-the-art PET/CT systems from different vendors and confirmed with reader studies, DDPET-3D consistently produced superior denoised results. Tested on images acquired at a different hospital, DDPET-3D also demonstrated its potential to be easily transferred to another medical center without further network fine-tuning. In previous literature, about 77% of deep-learning-based healthcare papers are limited to institutional or regionally distributed dataset [37]. Cross-center generalizability would be crucial to implement deep-learning algorithms in clinical practice.
Compared to previous diffusion models which are time-consuming to reconstruct the entire 3D volumes, once the DDPET-3D is fully-trained, the denoising process is efficient and can be finished within 15 min using one NVIDIA A40 GPU. This is sufficiently fast in reality as the clinical PET image reconstruction time is 15-35 min depending on reconstruction parameters [38] on the United Imaging uExplorer scanner.
Confirmed with reader studies by professional nuclear physicians, the proposed DDPET-3D can provide a similar or better image quality compared with normal-dose images in terms of noise suppression and structural fidelity. DDPET-3D effectively suppressed background noise and false-positive hot spots in certain patients, making true-positive lesion easier to identify.
[…]
In conclusion, the proposed DDPET-3D provides better or similar image quality in different low-dose settings compared to images acquired in a normal-dose setting. The experimental results showed that DDPET-3D achieved superior performance compared with previous methods. While this study focuses on PET image denoising, we believe that the proposed method could be easily extended to other 3D reconstruction tasks for different imaging modalities.
IV Methods
IV-A Diffusion models
@@
IV-B Conditional PET image denoising
@@
IV-C Proposed DDPET-3D
Figure 9: The training (top) and the sampling (bottom) pipeline of the proposed DDPET-3D method. The DDPET-3D uses multiple neighboring slices as additional inputs to predict the the central slice, allowing network to observe 3D information during training and testing. The DDPET-3D is also conditioned on the injected dose in order to accommodate inputs with varying noise levels. In sampling, we propose to fix the Gaussian latent for each slice in the 3D volume to address the inconsistency along z-axis. DDPET-3D also use a pre-trained denoiser prior to ensure accurate quantification in sampling.
@@ IV-C Proposed DDPET-3D
IV-D PET data acquisitions and reconstructions
@@
University Hospital of Bern Dataset: 377 subjects with the 18F-FDG tracer were included in this dataset. All data were acquired using a Siemens Biograph Vision Quadra total-body PET/CT system. Images were reconstructed using OSEM with 6 iterations and 5 subsets. A $5$\mathrm{m}\mathrm{m}$$ FWHM Gaussian filter was applied after reconstructions. The reconstruction matrix size was $644\times 440\times 440$ with a $1.65\times 1.65\times 1.65$\mathrm{m}\mathrm{m}$^{3}$ voxel size. We randomly selected 99 subjects for training, 10 subjects for validation, and 268 subjects for testing. The average full administered activity was $218.7\pm 49.4\ \mathrm{MBq}$ with scan duration of 10 min and post-injection tracer uptake time of 30 min. 1%, 2%, 5%, 10%, 25%, and 50% low-count levels were available through listmode rebinning.
Yale-New Haven Hospital Dataset: 195 subjects with the 18F-FDG tracer were included in this dataset. All data were acquired using a Siemens Biograph mCT PET/CT system. The whole-body scan protocol with continuous-bed motion scanning was used. Images were reconstructed using OSEM with 2 iterations and 21 subsets. A $5$\mathrm{m}\mathrm{m}$$ FWHM Gaussian filter was applied after reconstructions. The voxel size of the reconstructed image was $2.04\times 2.04\times 2.03$\mathrm{m}\mathrm{m}$^{3}$. We randomly selected 90 subjects for training, 10 subjects for validation, and 95 subjects for testing. The image size was $400\times 400$ in the transverse plane and varied in the axial direction depending on patient height. The average full administered activity was $256.3\pm 16.2\mathrm{MBq}$ with scan duration of 5 min every bed position and post-injection tracer uptake time of 60 min. 5%, 10%, and 20% low-count levels were available through listmode rebinning.
@@
V Ablation studies

### v2

# Dose-aware Diffusion Model for 3D Low-count PET: Multi-institutional Validation with Reader Study and Real Low-dose Data

@@ I Introduction
Positron Emission Tomography (PET) is a functional imaging modality widely used in oncology, cardiology, and neurology studies [1, 2, 3]. Given the growing concern of radiation exposure and potentially increased cancer risks accompanied with PET scans, reducing the PET injection dose is desirable [4]. In addition, reducing patient motion artifacts and patients’ discomfort through shorter scans as well as enhancing image quality on lower-performance scanners would also be beneficial in clinical settings. However, reduced injection dose and shorter scan duration can negatively affect PET image quality, and may affect diagnostic performance such as the identification of low-contrast lesions [5]. Therefore, reconstructing high-quality images from noisy input is an important topic. In fact, making the radiation dose as low as reasonably achievable (the ALARA principle) is a commonly accepted practice in clinical settings.
@@
Many different deep learning (DL) methods were proposed for low-dose/low-count PET image reconstructions/restorations [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. In particular, with the advancement of cutting-edge generative models [18, 19], diffusion models have demonstrated strong potential for low-dose PET imaging. Concretely, building upon the DDPM framework [20], Gong et al. proposed to perform PET image denoising with MRI as prior information for improved image quality [21]. Unfortunately, PET-MR systems are not widely available and it would be time-consuming to obtain sequential and paired PET-MR images in reality for network training and image reconstructions. Jiang et al.[22] adopted the latent diffusion model [23] for unsupervised PET denoising. Diffusion models have also been proposed for other imaging modalities, such as CT [24] and MRI [25, 26]. However, we observe that most of these methods focus on 2D imaging, leaving the application of diffusion models for 3D PET imaging largely underexplored in the literature. Several challenges should be addressed to effectively apply diffusion models to 3D PET imaging.
@@
First, PET imaging (PET) is intrinsically a 3D imaging modality, while previous 2D diffusion works fail to consider 3D spatial correlation between different slices. Thus, a 3D diffusion model is desirable. However, due to the hardware memory limit, directly extending the diffusion model to 3D would be difficult. There are a few previous works that aim to address the 3D imaging problems of diffusion models. For example, Chung et al.[27] proposed to apply a TV penalty term along the z-axis to remove inconsistencies within each reverse sampling step in the diffusion model. On the other hand, Lee et al.[28] utilized 2 pre-trained perpendicular 2D diffusion models to remove inconsistencies between slices. Nonetheless, presented later in this paper, these previous methods failed to produce satisfactory results for 3D low-count PET (e.g., inaccurate organ boundaries in extreme low-dose settings, inconsistent 3D reconstructions, distorted image features in some cases).
@@
Second, another challenge with PET image denoising is the high variation of image noise levels among patients, especially when dealing with multi-center, cross-scanner dataset. The noise level in PET images can be affected by various factors, including but not limited to: (1) Variations of post tracer injection acquisition start time. (2) Occasional tracer injection infiltration in some patients, causing the tracer to accumulate at the injection location, resulting in high image noise in the body. (3) The variation of patient’s BMI and associated weight-based tracer injection protocols. (4) Different hospitals using scanners with varying sensitivities have different total scan times for patients. (5) Variations of excretion into the urinary bladder. Higher excretion would also result in higher noise in other parts of the image. All these factors result in variability of noise levels in the reconstructed images. Therefore, a method to adaptively denoise images with varying noise levels is desirable. However, previously proposed methods mentioned above have limited generalizability to different noise levels. A network trained on one noise level often fails to produce high-quality reconstructions on other noise levels [16]. To address this problem, we previously proposed to combine multiple U-net-based [29] sub-networks with varying denoising power to generate optimal results for any input noise levels [16]. However, training multiple sub-networks requires tedious data pre-processing and long training time. The testing time also linearly increases with the number of sub-networks.
@@
Third, different from other imaging modalities, PET was developed as a quantitative tool, being recognized as providing an objective, and more accurate measure for prognosis and response monitoring purposes than visual inspection alone [30]. However, our experimental results showed that, although standard diffusion models (DDPM [20], DDIM[31]) produce visually appealing images, they typically fail to maintain accurate quantification and produced images with biased tracer uptake. Specifically, the tracer distribution in the entire image change after diffusion models. This issue is more severe in whole-body PET scans since the tracer uptake could vary significantly among organs.
@@
In this work, we developed a dose-aware diffusion model for 3D PET Imaging (DDPET-3D) to address these limitations. The main methodological contributions of the proposed DDPET-3D framework are: First: We proposed a 2.5D diffusion strategy with multiple fixed noise variables to address the 3D issue in diffusion models for medical imaging. DDPET-3D maintains a similar memory burden to 2D diffusion models while achieving high-quality reconstructions. Second: We proposed a dose-embedding strategy that allows noise-aware denoising. DDPET-3D can simultaneously denoise 3D PET images with varying low-dose/count levels. Third: To achieve a quantitative-accurate reconstruction, we proposed to use a denoised prior in DDPET-3D. The proposed denoised prior also allows the diffusion model to converge within 25 sampling steps. DDPET-3D can denoise 3D PET images within a reasonable time constraint ($\sim$15 mins on a single NVIDIA A40 GPU). Previous diffusion methods using DDPM sampling would require approximately $6$ hrs.
Furthermore, existing models are often tailored to specific hospitals, scanners, patient populations, acquisition protocols, and noise levels, limiting their generalizability across the diverse PET image acquisition settings encountered in practice. To address this, we evaluated the proposed method on a large-scale, cross-center, cross-scanner dataset, which is potentially the most diverse and extensive dataset in PET-related AI literature. In this work, we coordinated different hospitals around the world, prepared this large-scale dataset for the evaluation of PET denoising algorithms, and have made parts of the data publicly available. In addition to traditional image quality metrics (SSIM, PSNR, etc), we designed a reader study with nuclear medicine physicians to provide a clinical insight of the model performance. This evaluation highlights the method’s clinical potential and generalizability across various acquisition settings, making it more applicable to real-world scenarios.
@@
The ultimate goal of DL methods for PET image denoising is to achieve low-dose PET without compromising image quality and clinical performance. As outlined in Fig. 1(c), evaluated on real-world, large-scale, cross-center, ultra-low-dose 3D PET data, with confirmation from reader studies with nuclear medicine physicians, DDPET-3D demonstrates superior quantitative and qualitative results that are comparable or even better than the 100% full-count data as well as previous baseline methods across varying low-count-levels. DDPET-3D also showed superior generalizability when directly applied to data acquired at a different hospital without further network fine-tuning. We also evaluated the proposed method on a cohort of human studies with real low-dose injection. Presented results show the potential of DDPET-3D to achieve low-dose PET imaging while maintaining image quality. An overview of the dataset is provided in Fig. 1(b).
@@
II Methods
II-A Diffusion models
@@
II-B Conditional PET image denoising
@@
II-C Proposed DDPET-3D
Figure 2: The training (top) and the sampling (bottom) pipeline of the proposed DDPET-3D method. The DDPET-3D uses multiple neighboring slices as additional inputs to predict the central slice, allowing the network to observe 3D information during training and testing. The DDPET-3D is also conditioned on the injected dose in order to accommodate inputs with varying noise levels. In sampling, we propose to fix the Gaussian latent for each slice in the 3D volume to address the inconsistency along the z-axis. DDPET-3D also uses a pre-trained denoiser prior to ensure accurate quantification in sampling.
@@ II-C Proposed DDPET-3D
II-D PET data acquisitions and reconstructions
@@
University Hospital of Bern Dataset: 377 subjects with the 18F-FDG tracer were included in this dataset. All data were acquired using a Siemens Biograph Vision Quadra total-body PET/CT system. Images were reconstructed using OSEM with 6 iterations and 5 subsets. A $5$\mathrm{m}\mathrm{m}$$ FWHM Gaussian filter was applied after reconstructions. The reconstruction matrix size was $644\times 440\times 440$ with a $1.65\times 1.65\times 1.65$\mathrm{m}\mathrm{m}$^{3}$ voxel size. We randomly selected 99 subjects for training, 10 subjects for validation, and 268 subjects for testing. The average full administered activity was $218.7\pm 49.4\ \mathrm{MBq}$ with scan duration of 10 min and post-injection tracer uptake time of 60 min. 1%, 2%, 5%, 10%, 25%, and 50% low-count levels were available through listmode rebinning.
Yale-New Haven Hospital Dataset: 195 subjects with the 18F-FDG tracer were included in this dataset. All data were acquired using a Siemens Biograph mCT PET/CT system. The whole-body scan protocol with continuous bed motion scanning was used. Images were reconstructed using OSEM with 2 iterations and 21 subsets. A $5$\mathrm{m}\mathrm{m}$$ FWHM Gaussian filter was applied after reconstructions. The voxel size of the reconstructed image was $2.04\times 2.04\times 2.03$\mathrm{m}\mathrm{m}$^{3}$. We randomly selected 90 subjects for training, 10 subjects for validation, and 95 subjects for testing. The image size was $400\times 400$ in the transverse plane and varied in the axial direction depending on patient height. The average full administered activity was $256.3\pm 16.2\mathrm{MBq}$ with scan duration of 5 min every bed position and post-injection tracer uptake time of 60 min. 5%, 10%, and 20% low-count levels were available through listmode rebinning.
@@
III Results
III-A Reader study on three commercial scanners
Figure 4: Sample low-count denoised images reconstructed using the proposed DDPET-3D with different low-count levels. 6 different patients at different low-count levels were randomly chosen from the testing dataset. DDPET-3D can produce consistent denoised results for a wide range of different noise-levels. Ground-truth images were reconstructed using 100% full-count data. As determined using the 100% full-count images, blue arrows point to true lesions or hot spots in different patients, and green arrows point to false hot spots in the images. Blue dashed boxes indicate zoom-in ROI.
We performed a reader study to evaluate the network performance. 45 patient studies acquired using three different commercial scanners from three hospitals (15 from each hospital) were randomly selected from the testing dataset and included for the reader study. The three types of commercial scanners are the Siemens mCT scanner, the Siemens Vision Quadra scanner, and the United Imaging uExplorer scanner. Three hospitals are the Yale New Haven Hospital in the US, Shanghai Ruijin Hospital in China, and the University Hospital of Bern in Switzerland.
Low-count images were reconstructed by down-sampling the PET listmode events. For the images acquired on the Siemens Vision Quadra and the United Imaging uExplorer scanners, six low-count levels were available, including 1%, 2%, 5%, 10%, 25%, and 50%. For the images acquired using the Siemens mCT scanner, three low-count levels were available, including 5%, 10%, and 20%. Images were reconstructed using vendors’ software from United Imaging Healthcare and Siemens Healthineers. Details on image reconstruction are provided later in this paper (see Methods).
All three readers agreed that the proposed DDPET-3D can generate denoised results that are better or at least comparable to the 100% full-count images in terms of overall image quality across the three types of commercial scanners at different hospitals. Sample denoised images acquired at the Shanghai Ruijin Hospital (United Imaging uExplorer scanner) are presented in Fig. 4. DDPET-3D can generate consistent denoised results across a wide range of low-count levels. DDPET-3D was able to recover fine details in the images (e.g., rib bone in 1% results; myocardium, spine, and femur in 2% results). DDPET-3D also effectively suppressed background noise (e.g., liver in patients shown in Fig. 4), and lesions became clearer in the denoised images (blue arrows in Fig. 4). DDPET-3D effectively removed false hot spots (green arrows in the 25% results) in the images and the true lesions/hot spots are easier to identified. True lesions/hot spots and false hot spots were determined based on the 100% full-count images.
III-B Reader study on real low-dose scans
Figure 6: Sample real low-dose denoised images generated using the proposed DDPET-3D and the UNN[16] methods. Blue arrow points to the aorta wall, which was better reconstructed in DDPET-3D. No full-dose images are available for low-dose scans.
III-C Generalizability test in a different hospital
Figure 7: Low-count denoised images reconstructed using the proposed DDPET-3D with different low-count levels for data acquired at the UC Davis Medical Center. The model trained with Shanghai Ruijin data was directly applied for image reconstructions. This illustrates the DDPET-3D may be generalizable to different medical centers without further fine-tuning.
Ten patient studies from the UC Davis Medical Center were included in this experiment. Five low-count levels were generated through listmodel rebinning, including 2.5%, 6.25%, 12.5%, 25%, and 50%. Sample denoised results acquired at the UC Davis Medical Center are presented in Fig. 7. The proposed DDPET-3D showed superior generalizability when directly applied to data acquired at a different hospital. Corresponding quantitative measurements are included in Table I.
III-D Comparison with other baseline denoising methods
Figure 9: Comparison between DDPET-3D and other baseline denoising methods. Blue arrows point to a hot spot that was better recovered by the proposed DDPET-3D. Dashed blue line indicate the location of the transverse slice. All the images were displayed on the same scale and normalized by the total injected activities.
TABLE I: Quantitative assessment for different methods on all the testing patients from United Imaging and Siemens scanners. The measurements were obtained by averaging the values on the testing human studies. The proposed method consistently produced promising denoised results regardless of input count levels. The best results among different low-count levels are marked in red. Note that 674 out of 774 testing patients from the Shanghai Ruijin Hospital do not have 50% count reconstructions. $\uparrow$ and $\downarrow$ indicate that higher and lower values are better, respectively, for each metric.
[…]
Input 44.929 / 0.474 / 0.831 48.040 / 0.331 / 0.900 51.272 / 0.227 / 0.947
UNN 48.975 / 0.291 / 0.933 50.514 / 0.243 / 0.954 51.970 / 0.205 / 0.969
[…]
TABLE II: Quantitative assessment for different methods. The measurements were obtained by averaging the values on the testing human studies. The proposed method consistently produced promising denoised results regardless of input count levels. $p<0.01$ was observed in all groups when comparing the proposed method with all other methods (except for 50% SSIM values). The best results among different low-count levels are marked in red. Black voxels were removed for calculations. $\uparrow$ and $\downarrow$ indicate that higher and lower values are better, respectively, for each metric.
[…]
The proposed DDPET-3D was compared with UNN [16], DiffusionMBIR [27], and TPDM [28]. As described previously, UNN was chosen because it was also proposed to achieve noise-aware denoising and was one of the winning methods in 2022 Ultra Low-dose PET Imaging Challenge. DiffusionMBIR and TPDM were selected because they were recent works proposed to address the diffusion 3D inconsistencies problem in medical imaging. We also compared with the standard DDIM (Denoising Diffusion Implicit Models) sampling [31]. Detailed implementations of different methods will be provided later in this paper (see Methods).
As presented in Fig. 9, standard DDIM sampling produced images with severe inconsistencies between slices in the sagittal view. Both DiffusionMBIR and TPMD improved the results, but the problem still exists. Proposed DDPET-3D method produced noticeably more consistent reconstructions. Moreover, even though other diffusion models produced visually realistic PET images, they failed to accurately reconstruct the structure of different organs. For example, all other diffusion models produced images with distorted myocardium as shown in the transverse slice in Fig. 9. The proposed DDPET-3D method also recovered some fine details in the images, such as the hot spot pointed by the blue arrows in Fig. 9, which is either over-smoothed (in UNN) or barely visible (in other diffusion methods).
Images were quantitatively evaluated using SSIM (Structural Similarity Index), PSNR (Peak signal-to-noise ratio), and NRMSE (Normalized Root-mean-square error). To facilitate the testing process, we tested all the comparison methods using 20 patients from the entire testing dataset. There were 6 different low-count levels of all the patients, resulting in a total of 120 testing studies. Quantitative results are presented in Table II.
III-E Albation studies